# Constraining pseudo-Diracness with astrophysical neutrino flavors

### v1

@@ I Introduction
where $m_{j}^{2}$ are the standard mass squared eigenvalues and $\delta m_{j}^{2}$ are the three new mass squared splittings which have not been observed experimentally and are expected to be smaller than the solar mass splitting. Due to the pseudo-Dirac structure, the mixing between the SM and the right-handed neutrinos are close to maximal. Under the assumption of maximal mixing, strong experimental constraints using solar neutrino data have been derived: $\delta m_{1,2}^{2}\lesssim 10^{-11}\,\textrm{eV}^{2}$, while it is not sensitive to $\delta m_{3}^{2}$ due to small $\theta_{13}$ Cirelli et al. (2005); de Gouvea et al. (2009); Anamiati et al. (2018); Ansarifard and Farzan (2023). Using atmospheric, beam and reactor neutrino data which are sensitive to atmospheric mass splitting and $\theta_{13}$, a much weaker constraint is obtained $\delta m_{3}^{2}\lesssim 10^{-5}\,\textrm{eV}^{2}$ Anamiati et al. (2018). (See also Cirelli et al. (2005).) Due to the small mass splitting, astrophysical neutrinos coming from distance sources of Mpc to Gpc allow to probe mass squared difference much smaller than $10^{-12}\,\textrm{eV}^{2}$ Keranen et al. (2003); Beacom et al. (2004); Cirelli et al. (2005); Esmaili (2010); Esmaili and Farzan (2012); Joshipura et al. (2014); Brdar and Hansen (2019); De Gouvêa et al. (2020); Martinez-Soler et al. (2022); Rink and Sen (2024); Carloni et al. (2024); Dixit et al. (2024). While the pseudo-Diracness of neutrinos are interesting in itself, ref. Fong et al. (2021) has showed that it could be linked to the cosmic baryon asymmetry for pseudo-Dirac seesaw model where the new squared mass splitting is intimately connected to the CP violation $\epsilon$ required for baryogenesis through leptogenesis as follows [^2] The pseudo-Dirac leptogenesis model proposed in ref. Ahn et al. (2016) does not impose such a strong constraint, predicting only a lower bound on seesaw scale for a given pseudo-Dirac mass splitting.
@@
In this work, we will use astrophysical neutrino flavor ratios to probe pseudo-Dirac neutrinos Keranen et al. (2003); Beacom et al. (2004); Esmaili (2010); Brdar and Hansen (2019). In this approach, insensitivity to absolute neutrino flux is a double-edged sword. The advantage is that when there exist hierarchies between $\delta m_{i}^{2}$, stringent upper bounds on larger $\delta m_{i}^{2}$ can be derived as we will show in this work. The disadvantage is that if all $\delta m_{i}^{2}$ ($i=1,2,3$) are of the same order, we will lose the sensitivity. This is to be compared with studies which consider specific astrophysical sources where only certain ranges of $\delta m_{i}^{2}$ for the first few oscillations can be probed and, away from this region, the analysis is limited by absolute flux uncertainty De Gouvêa et al. (2020); Martinez-Soler et al. (2022); Rink and Sen (2024); Carloni et al. (2024); Dixit et al. (2024). If the flux from certain source can be determined quite accurately, and in particular, if all $\delta m_{i}^{2}$ are of the same order, this latter approach is superior. In the view that only very few astrophysical sources are identified, we will focus on flavor ratios approach and see how well this will fare. The paper is organized as follows: in Section II, we will review the parametrization of pseudo-Dirac model at low energy while in Section III, we will discuss the framework to calculate the oscillation probability in this scenario. Then, we will present our main results in Section IV and conclude in Section V.
@@ III.1 Neutrino flavor ratios
$P_{\alpha\beta}^{std}=\sum_{i=1}^{3}|U_{\alpha i}|^{2}|U_{\alpha i}|^{2}.$ (39)
@@ Acknowledgments
YP acknowledges the support by Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP) Contract No. 2023/10734-3. CSF acknowledges the support by FAPESP Contracts No. 2019/11197-6, 2022/00404-3, 2023/01467-1 and Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under Contract No. 304917/2023-0.

### v2

@@ I Introduction
where $m_{j}^{2}$ are the standard mass squared eigenvalues and $\delta m_{j}^{2}$ are the three new mass squared splittings which have not been observed experimentally and are expected to be smaller than the solar mass splitting. Due to the pseudo-Dirac structure, the mixing between the SM and the right-handed neutrinos are close to maximal. Under the assumption of maximal mixing, strong experimental constraints using solar neutrino data have been derived: $\delta m_{1,2}^{2}\lesssim 10^{-11}\,\textrm{eV}^{2}$, while it is not sensitive to $\delta m_{3}^{2}$ due to small $\theta_{13}$ Cirelli et al. (2005); de Gouvea et al. (2009); Anamiati et al. (2018); Ansarifard and Farzan (2023). (See also Franklin et al. (2023).) Using atmospheric, beam and reactor neutrino data which are sensitive to atmospheric mass splitting and $\theta_{13}$, a much weaker constraint is obtained $\delta m_{3}^{2}\lesssim 10^{-5}\,\textrm{eV}^{2}$ Anamiati et al. (2018). (See also Cirelli et al. (2005).) Due to the small mass splitting, astrophysical neutrinos coming from distance sources of Mpc to Gpc allow to probe mass squared difference much smaller than $10^{-12}\,\textrm{eV}^{2}$ Crocker et al. (2000, 2002); Cirelli et al. (2005); Esmaili (2010); Esmaili and Farzan (2012); Joshipura et al. (2014); Shoemaker and Murase (2016); Brdar and Hansen (2019); De Gouvêa et al. (2020); Martinez-Soler et al. (2022); Rink and Sen (2024); Carloni et al. (2024); Dixit et al. (2024). While the pseudo-Diracness of neutrinos are interesting in itself, ref. Fong et al. (2021) has showed that it could be linked to the cosmic baryon asymmetry for pseudo-Dirac seesaw model where the new squared mass splitting is intimately connected to the CP violation $\epsilon$ required for baryogenesis through leptogenesis as follows [^2] The pseudo-Dirac leptogenesis model proposed in ref. Ahn et al. (2016) does not impose such a strong constraint, predicting only a lower bound on seesaw scale for a given pseudo-Dirac mass splitting.
@@
In this work, we will use astrophysical neutrino flavor ratios to probe pseudo-Dirac neutrinos Crocker et al. (2000, 2002); Keranen et al. (2003); Beacom et al. (2004); Esmaili (2010); Shoemaker and Murase (2016); Brdar and Hansen (2019). In this approach, insensitivity to absolute neutrino flux is a double-edged sword. The advantage is that when there exist hierarchies between $\delta m_{i}^{2}$, stringent upper bounds on larger $\delta m_{i}^{2}$ can be derived as we will show in this work. The disadvantage is that if all $\delta m_{i}^{2}$ ($i=1,2,3$) are of the same order, we will lose the sensitivity. This is to be compared with studies which consider specific astrophysical sources where only certain ranges of $\delta m_{i}^{2}$ for the first few oscillations can be probed and, away from this region, the analysis is limited by absolute flux uncertainty De Gouvêa et al. (2020); Martinez-Soler et al. (2022); Rink and Sen (2024); Carloni et al. (2024); Dixit et al. (2024). If the flux from certain source can be determined quite accurately, and in particular, if all $\delta m_{i}^{2}$ are of the same order, this latter approach is superior. In the view that only very few astrophysical sources are identified, we will focus on flavor ratios approach and see how well this will fare. The paper is organized as follows: in Section II, we will review the parametrization of pseudo-Dirac model at low energy while in Section III, we will discuss the framework to calculate the oscillation probability in this scenario. Then, we will present our main results in Section IV and conclude in Section V.
@@ III.1 Neutrino flavor ratios
$P_{\alpha\beta}^{std}=\sum_{i=1}^{3}|U_{\alpha i}|^{2}|U_{\beta i}|^{2}.$ (39)
@@ Acknowledgments
YP acknowledges the support by Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP) Contract No. 2023/10734-3 and 2023/01467-1. CSF acknowledges the support by FAPESP Contracts No. 2019/11197-6, 2022/00404-3 and Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under Contract No. 304917/2023-0.
Note Added: A recent work Dev et al. (2024) has also considered effects of pseudo-Dirac neutrinos on astrophysical neutrino flavor ratios and showed that if there is an overdensity of cosmic neutrino background, matter effect can be relevant.